1. A 15 F capacitor is charged to 60 V and then disconnected from the battery. The plate
separation is then increased from 2.0 mm to 3.5 mm. calculate the amount of work
required to separate the plates. If the capacitor is submerged in distilled water, with a
dielectric constant of 80, determine;
a. the new capacitance, and
b. the potential difference across the capacitor.
1
Expert's answer
2019-08-14T09:47:35-0400

a.


"C'=C\epsilon \frac{d}{d'}=(15\cdot 10^{-6})(80) \frac{2}{3.5}=6.86\cdot 10^{-4}F"

b. From the conservation of charge:


"CU=C'U'"

"U'=U\frac{C}{C'}=60\frac{15\cdot 10^{-6}}{6.86\cdot 10^{-4}}=1.3\ V"